# Explainable Artificial Intelligence for Autonomous Driving: A Comprehensive Overview and Field Guide for Future Research Directions

### v4

@@ 5.8 Large Language Models and Vision-Language Models-based explanations
Figure 12: Live explanations for temporally evolving scenes by LINGO-1. Graphics adapted from the original source: [209]
consideration of human factors in the design and development of this technology and having effective human-machine alignment for trustworthy autonomous driving. In this sense, it is crucial that users-in-the-loop have some form of interaction with AVs during a journey. Motivated by this concept, some recent works have approached conversational user interface between people on board and AVs as a Video Question Answering (VideoQA) task [212, 213, 218]. Asking questions about the behavior of an autonomous system is a part of our intuition, and in the context of AVs, getting answers to traffic-related situations and autonomous car action-related questions can help users have a comfortable and reliable journey. Other practical applications of LLMs and VLMs are interpretable motion planning [215], chain-of-thought-based reasoning for control and decision-making [216, 219], action justification with a control signal [220], predicting the intent of other traffic actors via visual reasoning and cues [211], and understanding the role of video transformer based-explanations on safety of autonomous driving [222] (see Table LABEL:tab:llm_vlm_exp for more details on LLM and VLM-based explanations).
Overall, as an emerging AI technology, LLMs and VLMs have tremendously benefitted AVs from the interpretability aspect, as described in the above studies. However, it is also worthwhile to mention that there are still spaces for improvement of these models as fictitiously generated explanations (e.g., hallucinations) may have serious safety implications and high-stakes consequences for self-driving actions and human life. We describe such caveats and potential solutions in Subsection 7.5 as future work.
Based on a high-level overview of all these studies, driving explanations are generally multi-modal, context-dependent, and task-specific. Moreover, end-to-end learning has become even more popular for highly autonomous decision-making owing to the combination of powerful deep-learning approaches and the availability of advanced sensor devices. As self-driving decisions directly impact road users, the concept of explainability should be accompanied by the safety standards and principles established by transportation regulators. Motivated by this proposition, we can define explainable autonomous driving as follows:
Driven by this definition and the state-of-the-art works in the above sections, we present a general and conceptual XAI framework for end-to-end autonomous driving and show the necessary components and process steps to achieve regulatory-compliant next-generation AVs.
6 An XAI framework: integrating end-to-end control, regulatory compliance, and explanations
We propose a unifying framework in which methods for developing XAI, end-to-end learning, and safety components are combined to inform general processes of regulatory principles. Each of these components has a concrete role in our framework. In our recent study [15], we have covered a brief description of end-to-end learning for AVs. We extend the scope of that work and describe the essential elements of end-to-end autonomous driving, and the role of and potential challenges with explanations in such a setting. We describe these individual components as follows:
$E=\{e_{1},e_{2},...e_{n}\},$
and a compendium of actions
$A=\{a_{1},a_{2},...a_{n}\},$
$C:E\mapsto A.$
$e\in E$
to a relevant action
$a\in A.$
$srC=f(eeC,A).$
This requirement could be as pragmatic as some inspection of individual vehicle safety (for example, verifying basic safety functions of an individual vehicle for re-licensing). That said, this concept should be considered as a thorough compliance testing of $eeC$ components from vehicle manufacturers to certify their public safety under international and/or national transportation guidelines such as [49] and [57]. The general principles for acceptable functional safety of road vehicles are defined by the ISO 26262 standard [63]. According to this standard, there should be a safety certification development with evidence-based rationales: the vehicle should be able to meet the established functional safety requirement in its operational context. Part 6 of the ISO 26262 standard [223] is dedicated to end-product development for automotive applications within the software level. This guideline includes the design, development, testing, and verification of software systems in automotive applications. Based on these standards, there seem to be two fundamental approaches to confirming regulatory compliance, which we label confirmation of compliance by “simulation,” and confirmation of compliance by “verification,”. These steps are aligned with our observation regarding the role of XAI in confirming regulatory compliance. In the case of the process of establishing regulatory compliance by simulation, the idea is that a selected set of autonomous actions can be simulated, and then assessed to be satisfactory. This approach is perhaps the most familiar, as it arises naturally from an engineering development trajectory, where the accuracy of simulators determines the quality of compliance (e.g., [224]). The confidence of the established compliance is a function of the accuracy and coverage of the simulation. However, this compliance process can be very expensive and prone to safety gaps, especially when consensus on the properties and scope of a simulation is difficult to achieve. Thus, in general, the simulation part can be considered a ”driving school” for AVs: The designed and developed learning software system should be tested rigorously in this phase before such an autonomous system, as a holistic architecture, is deployed to a real vehicle in the physical environment and real roads.
@@ 1. An end-to-end control component: Given all possible instances of environment,
Moreover, through the extensive analysis of the reviewed studies, a holistic view suggests that automotive researchers and practitioners have investigated the safety and explainability aspects of AVs separately in general. Nevertheless, as autonomous driving is a safety-critical application of AI, these concepts must be explored in tandem, and explainability-guided safety improvement deserves more attention. While the majority of AVs’ explanations are post-hoc and provide an opportunity to improve the existing driving system via a retrospective analysis, there is also a need to develop intrinsically interpretable driving systems. One of the seminal studies on the latter topic is Rudin’s [235] work. The author analytically and empirically proves that instead of interpreting black-box ML models in a post-hoc manner, interpretable by-design models should be a preferred choice to design trusted autonomous systems. The author also justifies that the adopted trade-off concept between accuracy and model interpretability is not valid across all domains and domain-specific tasks, such as in computer vision problems. It is possible that an interpretable model can also lead to the same accuracy as the black-box model. From this perspective, we present potential XAI methods with examples and show the role of both live and post-hoc explanations for safe and trustworthy autonomous driving.
7 Mind the gap: Future perspective of XAI in autonomous driving
In this section, we propose potential XAI approaches that can bridge the gap between missing pieces in ongoing research on explainable AVs. We specifically show the current limitations and problems with modern autonomous driving and present guidelines to make AVs more transparent, safer, and robust.
7.1 Trustworthy and Debuggable Explainable Vision
7.1.1 Explainable vision through post-hoc explanations
Explainable vision-directed actions for AVs are based on how high-level features are used to detect and identify objects. The literature reviewed in Subsection 5.3 presents promising attempts to interpret CNN predictions. It is also important that these explanations help debug, find flaws, and improve the existing vision system. We show the importance of vision-based post-hoc explanations in autonomous driving with two examples. As an alternative to the Trolley problems, International Telecommunication Unit (ITU) recently initiated the “‘Molly problem” defined as “A young girl called Molly is crossing the road alone and is hit by an unoccupied self-driving vehicle. There are no eye- witnesses. What should happen next?” [236]. This is where the value of the post-hoc explanation strategy emerges: The vision system of an automated vehicle could provide a rationale on how it perceived, identified, and distinguished on-road objects and why it continued to drive, resulting in hitting the pedestrian. Such explanations are important for transparency in post-accident investigations under regulatory compliance. The second example is the hacking of Tesla’s Autopilot by the McAfee Advanced Threat Research team [237]. The team added a sticker to the label of the actual speed limit, 35 mph, and caused the car’s heads-up display to perceive it as 85 mph (see Figure 14). The car wrongly accelerated beyond the allowed limit in that traffic area.
Figure 14: Extending the middle of 3 by a sticker on the speed limit sign (left side) causes Tesla’s ADAS (right side) to read the limit as 85 mph [237]. The red circle on the right side has manually been added to show the causal effect of the performed hacking.
Even without a careful look at the modified speed limit, humans will not immediately understand why the car sped up in this example. While this is an intentional test performed by humans, similar confusion of the intelligent driving system may be caused by natural phenomena, such as adverse atmospheric conditions. For example, assume that the part of the speed limit sign is covered by muddy rain, and the vehicle’s ADAS perceives 3 as 8, increases speed, and potentially causes a traffic accident. So, we see the importance of a post-hoc explanation once again. “Speed limit shows 85 mph, accelerating” would be a simple, timely description to understand the reason for speeding up in case of a regulatory investigation. So, it turns out that the history of descriptive natural language generated along with each relevant action could be helpful to produce reliable post-hoc explanations for critical traffic scenarios. This technique is helpful to debug and improve the existing system as well.
7.1.2 Explainable vision through intrinsic and live explanations
As explained in Challange 3 in [235], developing interpretable models that concurrently provide explanations is a promising approach towards achieving a transparent vision system, particularly in object classification tasks. Initial specifications on this perspective use a concept of the prototype layer as an addition to a deep neural network: parts of a visual object are decomposed into pieces, and the prototype layer picks out some representative parts of the object during training. When given a new image during testing, the network tries to discover the similarities between those parts of the new image and the ones learned as prototypes in training. In this way, the deep network accumulates evidence from the prototypes and classifies the objects accordingly. Empirical studies of this perspective on classifying handwritten digits, cars, Fashion-MNIST dataset [238], and bird species and cars [239] report nearly the same accuracy as the original black-box models on which they were built. This simple technique inherently produces an explanation in parallel with a prediction of the neural network and does not require any further post-hoc explanation. In the context of autonomous driving, generating live explanations can contribute significantly to accident prevention. Assume that a self-driving vehicle has an in-vehicle person (i.e., a backup driver or a passenger). The vehicle provides a control (i.e., stop) button for emergency use. The built-in interface inside the vehicle shows there is no human ahead crossing the road and continues driving while there is a human ahead actually (i.e., a vision system malfunction). By noticing such an anomaly on time, the person on board can use the emergency button to slow down and/or stop the car and
Figure 15: Model-based RL vs. model-free RL from explainability perspective.
prevent the accident. This simple example shows that live explanations have potential use in autonomous driving and provide an opportunity for safe navigation of a vehicle.
7.2 Explainable Actions Through Model-based Reinforcement Learning
Characterizing AV actions as a sequential decision-making process along the trajectory, such exploration is typically performed within MDP. The most commonly explored approaches for AVs include three types of MDP: IL, model-free RL, and model-based RL, as reviewed in the studies above. IL intends to mimic the behavior of a human driver; this learning process is computationally expensive as it first must gather real-world driving data as training data [240]. Hence, driving policies obtained under this setting can not be controlled at test time [241]. Model-free RL algorithms learn by sensing the environment directly and do not have access to the transition dynamics (i.e., prior knowledge) of the explored environment [101]. Such exploration lacks the explainability of a learned policy: A model-free RL agent explores the driving environment without specific guidance if applied in autonomous driving. It may take a long time to learn an optimal driving policy. This problem is directly addressed in model-based RL: An advantage of model-based RL is that an agent learns the model of the environment first and then adjusts its learning policy according to the dynamics of the environment [242, 243, 244]. This kind of targeted exploration is typically called planning, which inherently makes the learning procedure explainable (see Figure 15). The idea of planning in RL is vital for proper decision-making and has been investigated in detail in the Dyna architectures [245, 246, 247]. The Dyna and its variation, the linear-Dyna architecture, concurrently learn the world model while learning the optimal policy through interacting with the world. Dyna’s planning process creates a predicted future trajectory from an initially provided imaginary state. Based on this structure, model projection generates an optimal action and simultaneously produces a predicted state and a predicted reward. Those last two components can then be visualized and analyzed and help us understand why the agent prefers selecting a particular action at a particular time step as a basis for explanation. As each sequential action of autonomous driving may need a causal explanation, the Dyna architecture and model-based RL, in general, can provide tremendous benefits with their explainability functionality.
7.3 Predictive Knowledge as Knowledge Representation
One of the essential steps for safe driving is to represent sensor-based knowledge accurately. For example, assume an autonomous vehicle approaches a four-way intersection. Based on the interaction with the operational surroundings, the intelligent driving system may take the “full stop,” “turn left,” “turn right,” or “go straight” actions. As prior knowledge, these four predictions can be considered as a set of possible actions that the agent (i.e., the car) may choose. This predictive approach to capturing domain knowledge is regarded as predictive knowledge in RL, and the idea has acquired growing attention in robotics and autonomous systems research [248, 249, 250, 251]. Moreover, such knowledge is also a reliable basis for explanations of the potential actions taken by an autonomous system. In this context, general value functions (GVFs) provide preliminary techniques for representing predictive knowledge. By definition, GVFs aim to capture long-term predictive summaries about actual observations of an RL agent [251, 252]. For instance, an RL agent’s actions for autonomous driving may be judged with questions, and GVFs can be used to represent the corresponding answers. Examples of questions are “What is the probability I will not face a red signal light in the next intersection?” or “What is the expected time till I arrive at my destination, given my current driving policy?” Preliminary experiments performed on robots have demonstrated GVFs’ representation value in learning, and their computational accuracy as a proof of concept has been established [251]. Moreover, Kearney et al. [253] have deeply investigated the role of GVFs from an explainable AI perspective and intuitively analyzed that the question and answer parameters of a GVF can be interpretable. Once an agent inspects its own experience and interprets the decisions to itself, it can also provide an explanatory basis to others through computational estimates. Recent applications of predictive knowledge, formalized as GVFs, on perception problems [254], and learning policies in the real-world autonomous driving [255] setting show the potential benefits of this concept for autonomous driving tasks. Hence, predictive knowledge computationally formalized as GVFs deserves more attention as a knowledge representation approach in autonomous driving.
7.4 Incorporating Commonsense Knowledge into Actions: Temporal Questions and Robust Question-driven Software Hierarchy
Another significant aspect of the intelligibility of an AV’s actions is strongly related to the proper design of its learning software system. This software system, as an end-to-end framework, must provide a rationale for each action taken at $t_{n}$ and explain how this particular action leads to the appropriate subsequent action in the $t_{n+1}$ time step. We can infer that hierarchical software, corresponding to such principles, is an appropriate structure to support explainable decision-making. Such a structure directly reflects our thoughts while driving. For instance, “Will the traffic light change from green to yellow soon?”, “Do the pedestrians ahead intend to cross the road?” or “Why is the car ahead changing its lane?” are just some representative questions that mirror our driving-related considerations while in motion. With this intuition, we can say that the hierarchical software system of an autonomous car can benefit from question-answering-based reasoning, such as using VQA and VideoQA, and users-in-the-loop can judge the temporal actions of AVs by asking meaningful questions to the conversational interface and receiving answers to their queries.
A critical challenge of a QA system is its robustness against tricky or adversarial questions. Traditionally, automotive user interfaces are designed to respond to users’ action-reflecting or conventional questions; however, humans can also ask deliberate questions to observe the interface’s response to the tricky questions. For instance, assume in a left turn scenario, a user enquires, “Why is the car turning to the right?” How should the dialogue system respond to such a question? It is also noteworthy to underscore that people can ask these questions unintentionally (i.e., consider people with visual impairment, color blindness, etc.), and the dialogue system must capture the semantics of such question and present a corrective answer, such as “ The car is actually turning left as traffic light allows it,” as a typical example. In this sense, adversarial VQA has been a recent area of interest to improve the robustness of QA models [256, 257]. However, this concept has not been explored well in autonomous driving and is a promising direction to enhance the robustness and faithfulness of automotive UX.
From the RL perspective, a suitable approach reflecting temporal states is the concept of options [258]. Options are a generalization of actions in which an RL agent has a policy for taking action with the terminal state. The well-established option-critic architecture is motivated by the concept of options [259]. That architecture can learn internal policies and terminal states of options and has proven effective in end-to-end learning of options in the Arcade Learning Environment (ALE). An inherent structure of the option-critic architecture makes it suitable for further development for the learning system of AVs. Driving-related questions are often temporal, and new questions can be generated for the subsequent actions just after a few seconds. The time sensitivity of driving decisions changes dynamically in real
Figure 16: Our approach to AV2.0 vs AV1.0, and potential advantages of AV2.0 over AV1.0 in terms of its AI software stack. The image of the autonomous vehicle has been taken from [260].
time and exposes the vehicle to different levels of risk. Naturally, actions with lower risks are preferred. Nevertheless, in terms of time and computation, we need to explore efficiently to assess the risk levels associated with the corresponding actions. It is possible that focusing only on increasing RL reward may not lead to desired actions in the long term. As an example, Zhang et al. [261] have shown that solely considering reward as a “gold” metric and ignoring levels of risk, as in traditional RL, is not always a perfect decision for an autonomous system, and an RL agent may fail to find the optimal policy with such an exploration. In contrast, incorporating different levels of risks with corresponding actions can help discover an optimal policy in environments through different transition and reward dynamics. Accordingly, a well-constructed question hierarchy and evaluation of risk levels for chosen actions can help achieve timely and informative explanations for AVs in critical traffic circumstances.
@@ 7.3 Predictive Knowledge as Knowledge Representation
7.5 Toward AV 2.0: Unifying vision, language, and action within Embodied AI for safe and explainable end-to-end autonomous driving
@@
While the above subsections describe the potential XAI approaches for vision and action modules individually, we also need to consider AVs’ learning software as a holistic driving system. Three decades of research, starting with ALVINN in 1988 [18] and further succeeding with the DARPA Grand Challenge [262], have achieved significant milestones with traditional AI software. However, recent breakthroughs in Foundation Models in terms of LLMs and VLMs motivate a transition to next-generation AVs. This generation of AVs has been referred to as AV2.0 by industry professionals [263, 264, 265]. The proposal is that the availability of integrated sensor suites, computational resources (i.e., GPU, TPU), and deep learning approaches can help AVs navigate via an end-to-end approach through adaptive learning, scaling, and generalization in complex driving environments. The ability to learn continually through interaction with the environment rather than relying on static datasets has resulted in the emergence of a new direction, labelled as “Embodied AI” [266, 267], and AV2.0 can move forward with such a learning approach. Effectively unifying vision, language, and action within Embodied AI can enable an autonomous car to navigate, interpret, and describe its high-level decisions in real-time. However, safety and explainability components of an end-end self-driving architecture must overcome fundamental challenges in AI described below:
@@
  • Avoiding negative side effects: Autonomous driving is primarily associated with the ability of a self-driving car to avoid accidents and maintain a safe distance from stationary and dynamic objects along the planned motion trajectory. However, the scope of the problem is not limited to this feature. Consider a scenario where an autonomous car interacts with another two vehicles, V1 and V2, at a specific moment. While aiming to make safe temporal decisions by itself, the autonomous car must also ensure that it does not implicitly enable V1 and V2 to cause an accident at that road segment as a part of vehicle-to-vehicle (V2V) communication. According to [268], one potential solution to this problem is to leverage cooperative Inverse RL [269], where an autonomous system can cooperate with humans, and a human actor can always shut down the autonomous system in case such a system exhibits undesirable behavior. In the context of autonomous driving, this nuance can be related to an AV’s communication with a human-operated vehicle or other remote operator monitoring an AV’s overall driving safety. One of the prominent methods in this context is Sympathetic Cooperative Driving or SymCoDrive paradigm [270], which trains agents not only to achieve safe driving for themselves but also for human-controlled vehicles by promoting altruistic driving behavior in cooperative autonomous driving. As the deployment of AVs on roads is a gradual process, synergy with human-operated vehicles is a viable approach for socially aware and safe navigation.
@@
  • Scalable oversight: Can humans measure whether AVs perform at a human level or better in general in all driving situations, where in specific moments, evaluating the driving behavior of an end-to-end AV performance may be difficult for humans due to some reasons (i.e., refer to the Molly problem above [236]). While being outside of human override, temporarily, for various reasons, can we trust that AVs will behave safely at that moment? Amodei et al. [268] report that a potential solution to this problem may be semi-supervised RL: an agent can see its reward on a small subset of episodes or times steps. While rewards from all episodes are used to evaluate the agent’s performance, the agent can only use that subset of rewards to optimize its performance under this setting.
@@
These problems reflect a broad spectrum of potential safety issues with end-to-end AVs. However, we argue that the proposal misses yet another essential concept, namely fail-safe ability. This concept has been investigated in some recent work [273, 274, 275]; however, the recent proposals of the next-generation AVs do not explicitly consider this functionality as an integral component of this technology. Human drivers often have a rest once they feel tired on long trips, and a short rest may help them feel mentally/physically better in the next phase of their driving. The same example can applied to AVs as well. Due to internal reasons (e.g., temporary system malfunction) or external factors (e.g., extremely adverse weather conditions), AVs may need to pause their trip temporarily and prevent further high-stakes consequences ahead. Such capability should not be considered a limitation of AVs; on the contrary, it is an optimal design strategy that foresees potential issues due to any factors and makes AVs behave safely by directing them to “have a short rest.”
@@
Explainability: The reviewed studies in Section 5 show a significant milestone in the explainability of self-driving systems. However, there are still significant gaps and challenges to achieving accurate and timely explanations in all phases of trips. For instance, as of September 2023, it is reported that LINGO-1 exhibits roughly 60% performance in its natural language and VQA-based explanations compared to human-level performance [209].
@@
Apart from informational content, another critical aspect deserving attention is the timing perspective of such explanations: the lead time for emergent scenarios, perhaps using extensive scenario-based evaluations or case-based reasoning, must be engineered appropriately. Furthermore, a well-known problem with large pre-trained models, hallucinations, is another challenge in explanation delivery. Particularly, in QA models, the model must generate a response based on the joint question and scene-based semantics rather than being influenced by the question itself, such as in the case of adversarial QAs described in the previous subsection. We have recently performed an empirical study [222] on the latter and shown that even advanced VLMs can fail to detect the language bias in QA models and present incorrect explanations in case of human adversarial questions. This issue, in turn, may damage user trust and can also have negative safety implications for self-driving. So, we argue that large pre-trained models’ construction mechanism can be adjusted and regulated with common sense and human-defined concepts, as also posited by Kenny and Shah [276]. Hence, designing robust QA models deserves more attention to enable meaningful and trustworthy dialogues between users and AVs. These features are key for achieving effective human-AI alignment [277, 278], trust [279, 280, 281], and public acceptance [67, 281, 282] with AV2.0. We describe our approach to AV2.0 and its difference from AV1.0 in Figure 16.

### v5

@@ 1. An end-to-end control component: Given all possible instances of environment,
Duan et al., [187], 2024 Unifying imitation learning with LLMs to enhance safety of end-to-end driving Vicuna LLM, Swin transformer Textual All groups
in this phase before such an autonomous system, as a holistic architecture, is deployed to a real vehicle in the physical environment and real roads.
@@
In the next section, we envision the future of AVs research in the realm of safety and explainability based on current industrial trends and list some potential challenges toward this goal.
@@
6 Toward AV 2.0: Unifying vision, language, and action within Embodied AI for safe and explainable end-to-end autonomous driving
@@
While the above subsections primarily describe the potential XAI approaches from the perspective of specific components, we also need to consider AVs’ learning software as a holistic driving system. Three decades of research, starting with ALVINN in 1988 [14] and further succeeding with the DARPA Grand Challenge [199], have achieved significant milestones with traditional AI software. However, recent breakthroughs in Foundation Models in terms of LLMs and VLMs motivate a transition to next-generation AVs. This generation of AVs has been referred to as AV2.0 by industry professionals [200, 201, 202]. The proposal is that the availability of integrated sensor suites, computational resources (i.e., GPU, TPU), and deep learning approaches can help AVs navigate via an end-to-end approach through adaptive learning, scaling, and generalization in complex driving environments. The ability to learn continually through interaction with the environment rather than relying on static datasets has resulted in the emergence of a new direction, labelled as “Embodied AI” [203, 204], and AV2.0 can move forward with such a learning approach. Effectively unifying vision, language, and action within Embodied AI can enable an autonomous car to navigate, interpret, and describe its high-level decisions in real-time. However, safety and explainability components of an end-end self-driving architecture must overcome fundamental challenges in AI described below:
@@
  • Avoiding negative side effects: Autonomous driving is primarily associated with the ability of a self-driving car to avoid accidents and maintain a safe distance from stationary and dynamic objects along the planned motion trajectory. However, the scope of the problem is not limited to this feature. Consider a scenario where an autonomous car interacts with another two vehicles, V1 and V2, at a specific moment. While aiming to make safe temporal decisions by itself, the autonomous car must also ensure that it does not implicitly enable V1 and V2 to cause an accident at that road segment as a part of vehicle-to-vehicle (V2V) communication. According to [205], a potential solution
    Figure 8: Our approach to AV2.0 vs AV1.0, and potential advantages of AV2.0 over AV1.0 in terms of its AI software stack, safety and explainability. The image of the vehicle has been taken from [206].
    to this problem could be to leverage cooperative Inverse RL [207], where an autonomous system can cooperate with humans, and a human actor can always shut down the autonomous system in case such a system exhibits undesirable behavior. In the context of autonomous driving, this nuance can be related to an AV’s communication with a human-operated vehicle or other remote operator monitoring an AV’s overall driving safety. One of the prominent methods in this context is Sympathetic Cooperative Driving or SymCoDrive paradigm [208], which trains agents not only to achieve safe driving for themselves but also for human-controlled vehicles by promoting altruistic driving behavior in cooperative autonomous driving. As the deployment of AVs on roads is a gradual process, synergy with human-operated vehicles is a viable approach for socially aware and safe navigation.
@@
  • Scalable oversight: Can humans measure whether AVs perform at a human level or better in general in all driving situations, where in specific moments, evaluating the driving behavior of end-to-end driving may be difficult for humans due to some reasons. While being outside of human override, temporarily (i.e., refer to the Molly problem [210]), for various reasons, can we trust that AVs will behave safely at that moment? Amodei et al. [205] report that a potential solution to this problem may be semi-supervised RL: an agent can see its reward on a small subset of episodes or times steps. While rewards from all episodes are used to evaluate the agent’s performance, the agent can only use that subset of rewards to optimize its performance under this setting.
@@
These problems reflect a broad spectrum of potential safety issues with end-to-end AVs. However, we argue that the proposal misses yet another essential concept, namely fail-safe ability. This concept has been investigated in some recent work [212, 213, 214]; however, the recent proposals of the next-generation AVs [200, 201, 202] do not explicitly consider this functionality as an integral component of this technology. Human drivers often have a rest once they feel tired on long trips, and a short rest may help them feel mentally/physically better in the next phase of their driving. The same example can applied to AVs as well. Due to internal reasons (e.g., temporary system malfunction) or external factors (e.g., extremely adverse weather conditions), AVs may need to pause their trip temporarily and prevent further high-stakes consequences ahead. Such capability should not be considered a limitation of AVs; on the contrary, it is an optimal design strategy that foresees potential issues due to any factors and makes AVs behave safely by directing them to “have a short rest.”
@@
Explainability: The reviewed studies in Section 4 show a significant milestone in the explainability of self-driving systems. However, there are still significant gaps and challenges to achieving accurate and timely explanations in all phases of trips. For instance, as of September 2023, it is reported that LINGO-1 exhibits roughly 60% performance in its linguistic and VQA-based explanations compared to human-level performance [168].
@@
Apart from informational content, another critical aspect deserving attention is the timing perspective of such explanations: the lead time for emergent scenarios, perhaps using extensive scenario-based evaluations or case-based reasoning, must be engineered appropriately. Furthermore, a well-known problem with large pre-trained models, hallucinations, is another challenge in explanation delivery. Particularly, in QA models, the model must generate a response based on the joint question and scene-based semantics rather than being influenced by the question itself, such as in the case of adversarial QAs. We have recently performed an empirical study [179] on the latter and shown that even advanced VLMs can fail to detect the language bias in QA models and present incorrect explanations in case of human adversarial questions. This issue, in turn, may damage user trust and can also have negative safety implications for self-driving. So, we argue that large pre-trained models’ construction mechanism can be adjusted and regulated with common sense and human-defined concepts, as also posited by Kenny and Shah [215]. Hence, designing robust QA models deserves more attention to enable meaningful and trustworthy dialogues between users and AVs. These features are key for achieving effective human-AI alignment [216, 217], trust [7, 218, 219], and public acceptance [45, 219, 220] with AV2.0. We describe our approach to AV2.0 and its difference from AV1.0 in Figure 8.